OpenStudy (anonymous):

Find the equation in standard form of the ellipse with foci (0,5), and (0,-5) and major axis of length 14. a.)x^2/49+y^2/24=1 b.)x^2/24+y^2/49=1 c.)x^2/4+y^2/49=1 d.)x^2/49+y^2/4=1

OpenStudy (anonymous):

foci (0,5) , (0,-5) we get C=5 major axis of length 14 that means 2b=14 b=7 \[\frac{ x^2 }{ b^2 }+\frac{ y^2 }{ a^2 }=1\] \[a>b>0\]

OpenStudy (anonymous):

srry. 2a=14 a=7 :)

OpenStudy (anonymous):

\[c^2=a^2-b^2\] \[b^2=a^2-c^2\] \[b^2=49-25=24\] \[\frac{ x^2 }{ 24 }+\frac{ y^2 }{ 49 }=1\]

OpenStudy (anonymous):

answer is b

OpenStudy (anonymous):